OpenStudy (mayankdevnani):

A 5 quintal car is moving with a velocity of 54 km/hr. What is its impulse if it is stopped within 0.5 s by application of backward force ? Also determine the force applied.

OpenStudy (mayankdevnani):

@ghazi @mathslover @experimentX @hartnn

OpenStudy (ghazi):

\[impulse= \frac{ \delta force }{ \delta time }\]

OpenStudy (mayankdevnani):

no , impulse = force * time

OpenStudy (ghazi):

find acceleration first

OpenStudy (ghazi):

sorry for that

OpenStudy (mayankdevnani):

ok! left it.. what are say

OpenStudy (ghazi):

V= 54km/hr= 54*5/18= 5*3m/s=15m/s

OpenStudy (ghazi):

now again V= at 15=a*0.5 a=15/0.5=30m/s^2 (retardation)

OpenStudy (mayankdevnani):

hey.... 18= ??? where you got

OpenStudy (ghazi):

now force= mass* a convert that mass into kilograms

OpenStudy (mayankdevnani):

but, 1 quintal = 100 kg

OpenStudy (ghazi):

and then impulse = force* time , that is the short duration for which a huge force acted :)

OpenStudy (mayankdevnani):

then , @ghazi

OpenStudy (ghazi):

force= 500*30=15000 N and impulse= 15000*0.5= 7500 Ns

OpenStudy (mayankdevnani):

thnx..... @ghazi

OpenStudy (ghazi):

:)

OpenStudy (mayankdevnani):

What is the relation between Impulse and Momentum.

OpenStudy (mayankdevnani):

an it be @ghazi The Impulse is the change in momentum! Momentum (Kg m/s) = mass x velocity Impulase can also be worked out by the formula... Impulse (Ns) = Force x Time

OpenStudy (mayankdevnani):

is that right?????? @ghazi

OpenStudy (ghazi):

see momentum by newton's second law momentum= force/ time and impulse = force *time and you can see that higher the change in momentum lesser will be impulse basically impulse is inversely proportional to momentum

OpenStudy (ghazi):

clear?

OpenStudy (mayankdevnani):

ok clear.